When Mendel conducted his famous genetics experiments with peas, one sample of offspring consisted of 438 green peas and 152 yellow peas. Find the 90% confidence interval estimate for the percentage of yellow peas.
I tried this: 438+152= 590
152/590= .2576 1-.2576= .7424
90% confidence interval est. of yellow peas ????

Answer by stanbon(75887)

When Mendel conducted his famous genetics experiments with peas, one sample of offspring consisted of 438 green peas and 152 yellow peas. Find the 90% confidence interval estimate for the percentage of yellow peas.
---------
sample proportion = 152/(438+152) = 0.2576
---
standard error = invNorm(0.05)*sqrt(0.2576*0.7424/(438+152)) = 0.0296
----------------
90% CI: 0.2576 - 0.0296 < p < 0.2576 + 0.0296
90% CI: 0.2280 < p < 0.2872
=================================
Cheers,
Stan H.